# Self-Ensembling Gaussian Splatting for Few-Shot Novel View Synthesis

### v2

@@ 3.3 Self-Ensembling in 3DGS
Given the temporal samples around the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these samples. Concretely, we train a separate 3DGS model, named the $\mathbf{\Sigma}$-model, on the posed images without perturbations. To prevent the $\mathbf{\Sigma}$-model from being trapped at a suboptimal solution, we guide the training of the $\mathbf{\Sigma}$-model with a regularization [47] formulated as
@@
where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term [21], and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively. We also utilize a co-pruning strategy [47] to enhance the regularization. Therefore, the $\mathbf{\Sigma}$-model aggregates information from different samples in the parameter space, serving as an ensemble of 3DGS models. The final loss function during training is defined as
@@
where $\gamma=1$ by default and $\mathcal{L}_{\text{RGB}}$ represents a photometric loss over the training views. During testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis. Notably, compared with the previous approach [47], our SE-GS inherently encodes a large number of 3DGS models without significant computational overhead, thereby enabling effective regularization efficiently. Furthermore, the regularization is carried out on pseudo views in a self-supervised manner, making it independent of additional information such as depth [24, 53].
@@
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
Mip-NeRF [2] 16.11 22.91 24.88 0.401 0.756 0.826 0.460 0.213 0.160 0.215 0.090 0.066
DietNeRF [19] 14.94 21.75 24.28 0.370 0.717 0.801 0.496 0.248 0.183 0.240 0.105 0.073
RegNeRF [30] 19.08 23.10 24.86 0.587 0.760 0.820 0.336 0.206 0.161 0.149 0.086 0.067
FreeNeRF [44] 19.63 23.73 25.13 0.612 0.779 0.827 0.308 0.195 0.160 0.134 0.075 0.064
SparseNeRF [38] 19.86 - - 0.624 - - 0.328 - - 0.127 - -
3DGS [21] 19.22 23.80 25.44 0.649 0.814 0.860 0.229 0.125 0.096 0.120 0.066 0.051
DNGaussian [24] 19.12 - - 0.591 - - 0.294 - - 0.132 - -
FSGS [53] 20.43 24.09 25.31 0.682 0.823 0.860 0.248 0.145 0.122 0.104 0.066 0.054
CoR-GS [47] 20.45 24.49 26.06 0.712 0.837 0.874 0.196 0.115 0.089 0.101 0.060 0.046
SE-GS 20.79 24.78 26.36 0.724 0.839 0.878 0.183 0.110 0.084 0.100 0.059 0.045
@@
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
3DGS [21] 17.67 23.69 26.80 0.804 0.894 0.941 0.158 0.086 0.050 0.112 0.058 0.031
MVSplat [7] 17.33 16.34 16.22 0.598 0.596 0.587 0.279 0.296 0.304 - - -
DNGaussian [24] 18.57 - - 0.776 - - 0.178 - - 0.110 - -
FSGS [53] 17.84 23.68 26.17 0.822 0.905 0.941 0.161 0.096 0.064 0.110 0.056 0.036
CoR-GS [47] 18.65 24.39 27.38 0.835 0.910 0.950 0.140 0.074 0.045 0.099 0.049 0.028
SE-GS 19.24 25.28 28.08 0.857 0.924 0.958 0.132 0.073 0.043 0.089 0.045 0.026
@@ 4.1 Setup
Datasets. We conduct experiments on four datasets, i.e., LLFF [27], DTU [20], Mip-NeRF360 [3], and MVImgNet [45]. On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in [47], using the same training/testing splits and downsampling rates. As suggested in [47], we mask the background when assessing the quality of novel view synthesis on the DTU dataset. Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiments to a large-scale dataset, i.e., MVImgNet. We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512. The Gaussian parameters for all evaluated methods are initialized based on the point clouds obtained from COLMAP [36] to ensure a fair comparison. Note that COLMAP fails on certain DTU scenes in the sparse-view setting. In these cases, we instead randomly initialize the point cloud. Please refer to the supplementary material for more details on the setup.
@@ 4.2 Quantitative Results
We provide the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively. The best, second-best, and third-best results in each column are highlighted in red, orange, and yellow, respectively. For MVSplat [7], we employ the pretrained model released by the authors and evaluate it on DTU. Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantage of scene-level radiance field modeling for high-quality novel view synthesis. For these scene-level approaches, the performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis. In this context, our method surpasses all competitors across the four datasets. Note that in the sparse-view scenario, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
@@
Specifically, on the LLFF and DTU datasets, we train the methods using 3, 6, and 9 views. Our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF. The improvement is larger on DTU, reaching a gain of $0.59$. Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in [47]. On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS and AVGE, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations. Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset. As reported in Table 4, our method yields consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views. Our SE-GS also demonstrates better compatibility with relatively dense views. For instance, the PSNR improvement increases from $0.26$ to $0.59$ as the number of training views varies from 5 to 10.
@@
The methods that leverage auxiliary data terms, such as DNGaussian [24] and FSGS [53], sometimes outperform 3DGS. However, in some cases, such as with 9 views on LLFF, the PSNR of FSGS becomes worse than 3DGS. Since the data acquired through off-the-shelf approaches might be unreliable, the quality of the NVS results is consequently affected. Compared with FSGS and DNGaussian, CoR-GS [47] exhibits better stability, outperforming 3DGS in more scenarios. This finding highlights the promising potential of regularization in the sparse-view setting. However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset. By contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability. Our method is capable of avoiding specific local optima as the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
@@ 4.3 Qualitative Results
Figure 5: Qualitative results. The methods are trained on sparse views and the renderings of novel views are illustrated. The images are selected from the DTU and MVImgNet datasets.
Fig. 5 presents qualitative comparisons of 3DGS, CoR-GS, and our SE-GS, showing renderings from novel views. The models are trained on the DTU and MVImgNet datasets with 3 and 5 posed images, respectively. Our SE-GS produces fewer visual artifacts than the other methods, achieving better robustness against the challenge of few-shot training views. Moreover, our method captures finer details, particularly in areas with complex and repeated textures, such as the pineapple and watch strap. This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under sparse-view conditions.
@@ 4.4 Analysis
To shed more light on the effectiveness and efficiency of our SE-GS, we perform a comprehensive analysis of the introduced self-ensembling mechanism. The analysis starts by comparing our approach with CoR-GS trained with varying numbers of Gaussian models. The detailed results on LLFF with 3 training views are shown in Fig. 6. As shown by Fig. 6(a), the PNSR of CoR-GS increases as more Gaussian models are trained. Since adding more Gaussian models enhances the regularization effect in CoR-GS, this observation highlights the potential of improving 3DGS in the sparse-view setting by strengthening the regularization process. The PSNR reaches a peak with 6 Gaussian models, indicating an upper bound of CoR-GS. Our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in more diverse samples of Gaussian models compared to CoR-GS. This leads to more effective regularization, as reflected by the better PSNR score.
Moreover, Fig. 6(b) shows the training speed measured as the number of training iterations completed per second. The speed of CoR-GS significantly drops as the number of Gaussian models increases. This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models. By contrast, in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model. The temporal samples are generated through perturbations. We achieve a comparable training speed to CoR-GS with 2 models, showing that the perturbation process incurs a negligible additional cost. Note that only the $\mathbf{\Sigma}$-model is retained for NVS during inference, so the testing speed of SE-GS is the same as that of the original 3DGS. Consequently, our method is capable of efficiently enhancing the performance of 3DGS with sparse training views.
@@
Table 5: Effectiveness of the uncertainty-aware perturbations. Random indicates that we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model. Gradient-Aware denotes an alternative in which we perturb the Gaussians based on gradients. All methods are trained on LLFF with 3 views for 2000 iterations and the PSNR is utilized as the metric.
Finally, we assess the effectiveness of our uncertainty-aware perturbation strategy, comparing it with two alternatives. Specifically, as introduced in Sec. 3, a straightforward method for perturbing the $\mathbf{\Delta}$-model is to add random noise to the parameters of all Gaussians in the model. We denote this approach as Random. Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing the gradients. Specifically, we replace the indicator function in Eq. 11 with one based on gradient magnitudes. Gaussians with gradient magnitudes exceeding a threshold are perturbed. In practice, we adopt the same threshold as used in density control. We train these two alternatives and our approaches on LLFF with 3 views for 2000 iterations, which challenges the methods in terms of convergence speed. As listed in Table 5, Random performs the worst since it leads to many invalid Gaussian samples, negatively impacting convergence speed. Note that our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method can only utilize the gradients at the current training step. The improvement in PSNR shown in Table 5 demonstrates the superior effectiveness of our approach.
@@ 5 Conclusion
In this paper, we have presented a new self-ensembling mechanism that enhances the novel view synthesis of 3DGS with sparse training images. We have tackled the challenge of overfitting by training an ensemble named $\mathbf{\Sigma}$-model with the guidance of diverse temporal models sampled from the Gaussian parameter space. To obtain such samples, we have introduced an effective strategy that perturbs a $\mathbf{\Delta}$-model based on uncertainties computed from dynamically updated buffers of pseudo-view renderings. We have conducted experiments on LLFF, DTU, Mip-NeRF360, and MVImgNet, as well as a comprehensive analysis of our approach. The experimental results have demonstrated the effectiveness, stability, and efficiency of our SE-GS. In future work, we plan to incorporate an outlier identifier into our perturbation mechanism to reduce the impact of unreliable samples in regularization and facilitate the self-ensembling process.
@@
In our experiments, we perturb the $\mathbf{\Delta}$-model every 500 iterations by default. To evaluate the effect of different intervals, we test our method on the LLFF dataset with 3 training views, using intervals ranging from 100 to 900. As shown in Fig. 7, the curves represent the PSNR values obtained on testing views at different training iterations. Perturbing the $\mathbf{\Delta}$-model too frequently, e.g., with an interval of 100, significantly degrades performance, indicating a negative impact on our method. Conversely, large intervals, such as 900, reduce the number of temporal samples, thereby weakening the self-ensembling effect and leading to decreased performance compared to the default setting.
Moreover, we analyze the effect of the noise level by conducting experiments with varying values of $\sigma$ in Eq. 16. Specifically, we adjust the value of $\omega$ from 0.01 to 0.3, corresponding to increasing noise levels. We train our method with each specific noise level on the LLFF dataset using 3 training views. The results on testing views at different training iterations are shown in Fig. 8. Strong perturbations, such as those with $\omega$ values of 0.3 and 0.2, result in significantly worse performance compared to other settings. In these cases, the temporal models are sampled too far from the $\mathbf{\Delta}$-model in the Gaussian parameter space, reducing the consistency among temporal samples. The performance is also limited when a small $\omega$, such as 0.01, is used. In this scenario, all temporal models closely resemble the $\mathbf{\Delta}$-model, diminishing the benefits of self-ensembling. In contrast to using a fixed $\omega$, we adopt a decay function, where $\omega$ dynamically varies from 0.08 to 0.02 during training. This strategy achieves a good trade-off between the consistency and diversity of the temporal samples.
@@
To better understand the presented perturbation mechanism, in Fig. 9, we illustrate the percentage of perturbed Gaussians in the $\mathbf{\Delta}$-model. Note that the threshold $\tau$ in Eq. 15 is defined adaptively. Therefore, the percentage dynamically varies throughout training. This dynamic behavior balances the diversity and reliability of the temporal models, thereby enhancing the self-ensembling process.
@@
We provide more visualizations including pseudo-view renderings of the $\mathbf{\Sigma}$-model and the $\mathbf{\Delta}$-model, uncertainty maps, and NVS results. For more details, please refer to the uploaded videos. Notably, the videos of pseudo-view renderings and uncertainty maps showcase the results across different training iterations.

### v3

@@ 3.4 Self-Ensembling in 3DGS
Given the perturbed models derived from the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these models. To ensure efficiency, we perform self-ensembling during training. Concretely, we train a separate 3DGS model, named the $\mathbf{\Sigma}$-model, on the training views without perturbation. Its training is guided with an additional regularization formulated as
@@
where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term [21], and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and the perturbed model, respectively. We also utilize a co-pruning strategy [48] to enhance the regularization. Therefore, the $\mathbf{\Sigma}$-model aggregates information from diverse perturbed models, functioning as an ensemble of 3DGS models. The final loss function during training is defined as
@@
where $\gamma=1$ by default and $\mathcal{L}_{\text{RGB}}$ represents a photometric loss over the training views. During testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis. Notably, compared with the previous approach [48], our SE-GS inherently encodes diverse 3DGS models without significant computational overhead, thereby enabling efficient and effective regularization. Moreover, the regularization is applied on pseudo views in a self-supervised manner, making it independent of additional information such as depth [25, 54].
@@
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
Mip-NeRF [2] 16.11 22.91 24.88 0.401 0.756 0.826 0.460 0.213 0.160
DietNeRF [19] 14.94 21.75 24.28 0.370 0.717 0.801 0.496 0.248 0.183
RegNeRF [31] 19.08 23.10 24.86 0.587 0.760 0.820 0.336 0.206 0.161
FreeNeRF [45] 19.63 23.73 25.13 0.612 0.779 0.827 0.308 0.195 0.160
SparseNeRF [39] 19.86 - - 0.624 - - 0.328 - -
3DGS [21] 19.22 23.80 25.44 0.649 0.814 0.860 0.229 0.125 0.096
DNGaussian [25] 19.12 22.01 22.62 0.591 0.717 0.741 0.294 0.246 0.244
FSGS [54] 20.43 24.09 25.31 0.682 0.823 0.860 0.248 0.145 0.122
CoR-GS [48] 20.45 24.49 26.06 0.712 0.837 0.874 0.196 0.115 0.089
SE-GS 20.79 24.78 26.36 0.724 0.839 0.878 0.183 0.110 0.084
@@
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
3DGS [21] 17.67 23.69 26.80 0.804 0.894 0.941 0.158 0.086 0.050
MVSplat [7] 17.33 16.34 16.22 0.598 0.596 0.587 0.279 0.296 0.304
DNGaussian [25] 18.57 22.56 25.25 0.776 0.862 0.917 0.178 0.114 0.077
FSGS [54] 17.84 23.68 26.17 0.822 0.905 0.941 0.161 0.096 0.064
CoR-GS [48] 18.65 24.39 27.38 0.835 0.910 0.950 0.140 0.074 0.045
SE-GS 19.24 25.28 28.08 0.857 0.924 0.958 0.132 0.073 0.043
@@ 4.1 Setup
Datasets. We conduct experiments on four datasets, i.e., LLFF [28], DTU [20], Mip-NeRF360 [3], and MVImgNet [46]. On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in [48], using the same training/testing splits. As suggested in [48], we mask the background when assessing the quality of novel view synthesis on the DTU dataset. Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiments to a large-scale dataset, i.e., MVImgNet. We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512. Notably, in our experiments, the Gaussian parameters for all evaluated methods are initialized based on the same point clouds obtained from COLMAP [37], ensuring a fair comparison. Moreover, COLMAP fails on certain DTU scenes in the sparse-view setting. In these cases, we instead randomly initialize the point cloud. Please refer to the supplementary material for more details on the setup.
@@ 4.2 Quantitative Results
We provide the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively. The best, second-best, and third-best results in each column are highlighted in red, orange, and yellow, respectively. For MVSplat [7], we employ the pretrained model released by the authors and evaluate it on DTU. Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantage of scene-level radiance field modeling for high-quality novel view synthesis. For these scene-level approaches, the performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis. In this context, our method surpasses all competitors across the four datasets. Note that in sparse-view scenarios, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
@@
Specifically, our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF. The improvement is larger on DTU, reaching a gain of $0.59$. Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in [48]. On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations. Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset. As reported in Table 4, our method yields consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views. Our SE-GS also demonstrates better compatibility with relatively dense views. For instance, the PSNR improvement increases from $0.26$ to $0.59$ as the number of training views varies from 5 to 10.
@@
The methods that leverage auxiliary data terms, such as DNGaussian [25] and FSGS [54], sometimes outperform 3DGS. However, in some cases, such as with 9 views on LLFF, the PSNR of FSGS becomes worse than 3DGS. Since the data acquired through off-the-shelf approaches might be unreliable, the quality of the NVS results is consequently affected. Compared with FSGS and DNGaussian, CoR-GS [48] exhibits better stability, outperforming 3DGS in more scenarios. This finding highlights the promising potential of regularization in the sparse-view setting. However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset. In contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability. This is attributed to its ability to effectively aggregate information from diverse perturbed models in the self-ensembling paradigm, leading to robust regularization.
@@ 4.3 Qualitative Results
Figure 5: Qualitative results. The methods are trained on sparse views and the renderings of novel views are illustrated. The images are from the DTU and MVImgNet datasets.
Fig. 5 presents qualitative comparisons of 3DGS, CoR-GS, and our SE-GS, showing renderings from novel views. The models are trained on the DTU and MVImgNet datasets with 3 and 5 posed images, respectively. Our SE-GS produces fewer visual artifacts than the other methods, achieving better robustness against the challenge of few-shot training views. Moreover, our method captures finer details, particularly in areas with complex and repeated textures. This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under sparse-view conditions. Please refer to the supplementary material for more visualizations.
@@ 4.4 Analysis
To shed more light on the effectiveness and efficiency of our SE-GS, we perform a comprehensive analysis of the introduced self-ensembling mechanism. The analysis starts by comparing our approach with CoR-GS trained with varying numbers of Gaussian models. The detailed results on LLFF with 3 training views are shown in Fig. 6. As shown by Fig. 6(a), the PNSR of CoR-GS increases as more Gaussian models are trained. Since adding more Gaussian models enhances the regularization effect in CoR-GS, this observation highlights the potential of improving 3DGS in the sparse-view setting by strengthening the regularization process. The PSNR reaches a peak with 6 Gaussian models, indicating an upper bound of CoR-GS. Our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in more diverse 3DGS models compared to CoR-GS. This leads to more effective regularization, as reflected by the better PSNR score.
Moreover, Fig. 6(b) shows the training speed measured as the number of training iterations completed per second. The speed of CoR-GS significantly drops as the number of Gaussian models increases. This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models. In contrast, in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model. The perturbed models are generated through perturbation. In this context, we achieve a comparable training speed to CoR-GS with 2 models, showing that the perturbation process incurs a negligible additional cost. Note that only the $\mathbf{\Sigma}$-model is retained for NVS during inference, so the testing speed of SE-GS is the same as that of the original 3DGS. Consequently, our method is capable of efficiently enhancing the performance of 3DGS with sparse training views.
@@
Table 5: Effectiveness of the uncertainty-aware perturbations. Random indicates that we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model. Gradient-Aware denotes an alternative in which we perturb the Gaussians based on gradients. All methods are trained on LLFF with 3 views for 2000 iterations, and PSNR is utilized as the metric.
Figure 7: Impact of model diversity. The $x$-axis represents model agreement, indicating the similarity among models involved during training. Higher agreement values suggest greater model similarity and lower diversity. The $y$-axis denotes the PSNR of the methods during testing. The experiment is conducted on LLFF with three training views.
Finally, we assess the effectiveness of our uncertainty-aware perturbation strategy, comparing it with two alternatives. As introduced in Sec. 3, a straightforward method for perturbing the $\mathbf{\Delta}$-model is to add random noise to the parameters of all Gaussians. We denote this approach as Random. Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing the gradients. Specifically, we replace the indicator function in Eq. 11 with one based on gradient magnitudes. Gaussians with gradient magnitudes exceeding a threshold are perturbed. In practice, we adopt the same threshold as used in density control. We train these two alternatives and our approaches on LLFF with 3 views for 2000 iterations, which challenges the methods in terms of convergence speed. As listed in Table 5, Random performs the worst, indicating that this strategy negatively impacts convergence speed. Notably, our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method only utilizes the gradients at the current training step. The effectiveness of this design is evidenced by the improvement in PSNR shown in Table 5. Additionally, to further investigate the impact of model diversity, we conduct a detailed analysis on LLFF. As shown in Fig. 7, the $x$-axis denotes the model agreement measured as PSNR between renderings of all involved models on the test views; the $y$-axis shows PSNR between the NVS result and the ground-truth image during testing. The Model agreement reflects consistency, where higher values indicate greater similarity among models. The agreements of CoR-GS with 5 and 8 models are nearly identical, meaning that the additional models lack sufficient diversity to enhance the ensembling process. This aligns with the observation in Fig. 6(a) that merely using more models yields limited improvement. While perturbation can enhance model diversity, it may weaken supervision reliability. As shown by the red triangle in Fig. 7, adding noise to all Gaussians yields overly diverse models. The supervision in the regularization term becomes too noisy, resulting in limited PSNR. In contrast, our uncertainty-aware perturbation method only perturbs unreliable Gaussians, enabling a good trade-off between model diversity and supervision reliability.
@@ 5 Conclusion
In this paper, we have presented a new self-ensembling mechanism that enhances the novel view synthesis of 3DGS with sparse training images. We have tackled the challenge of overfitting by training an ensemble named $\mathbf{\Sigma}$-model with the guidance of diverse perturbed models derived from a $\mathbf{\Delta}$-model. To obtain such models, we have introduced an effective strategy that perturbs the $\mathbf{\Delta}$-model based on uncertainties computed from dynamically updated buffers of pseudo-view renderings. We have conducted experiments on LLFF, DTU, Mip-NeRF360, and MVImgNet, as well as a comprehensive analysis of our approach. The experimental results have demonstrated the effectiveness, stability, and efficiency of our SE-GS. In future work, we plan to incorporate an outlier identifier into our perturbation mechanism to reduce the impact of unreliable models in regularization and facilitate the self-ensembling process.
@@
Details on Ensembling Learning
Figure 8: Pipeline of self-ensembling. Network variants are generated via dropout, and self-ensembling is achieved by utilizing a regularization term between the variants.
Ensemble learning has been recognized for its effectiveness in improving robustness and generalization [13], especially in scenarios with limited training data. We illustrate the training pipeline of existing self-ensembling approaches [11] in Fig. 8. Specifically, self-ensembling aims to aggregate information of diverse models without training multiple models from scratch. The model diversity is typically enhanced via dropout. In this context, the variants are dynamically derived from a single network during training, thereby prompting both diversity and efficiency. Self-ensembling is then achieved by incorporating a regularization term, which operates alongside the task-specific loss function based on ground truth. This regularization enforces consistency among the network variants by constraining their outputs, thereby mitigating model bias and enhancing robustness and generalization. Inspired by its effectiveness, we introduce the first self-ensembling Gaussian Splatting method to improve novel view synthesis in scenarios with sparse training views.
@@
In our experiments, we perturb the $\mathbf{\Delta}$-model every 500 iterations by default. To evaluate the effect of different intervals, we test our method on the LLFF dataset with 3 training views, using intervals ranging from 100 to 900. As shown in Fig. 9, the curves represent the PSNR values obtained on testing views at different training iterations. Perturbing the $\mathbf{\Delta}$-model too frequently, e.g., with an interval of 100, significantly degrades performance, indicating a negative impact on our method. Conversely, large intervals, such as 900, reduce the number of perturbed models, thereby weakening the self-ensembling effect and leading to decreased performance compared to the default setting.
Moreover, we analyze the effect of the noise level by conducting experiments with varying values of $\sigma$ in Eq. 16. Specifically, we adjust the value of $\omega$ from 0.01 to 0.3, corresponding to increasing noise levels. We train our method with each specific noise level on the LLFF dataset using 3 training views. The results on testing views at different training iterations are shown in Fig. 10. Strong perturbation, such as those with $\omega$ values of 0.3 and 0.2, result in significantly worse performance compared to other settings. In these cases, the perturbed models are too far from the $\mathbf{\Delta}$-model in the Gaussian parameter space, reducing the consistency among the variants. The performance is also limited when a small $\omega$, such as 0.01, is used. In this scenario, all perturbed models closely resemble the $\mathbf{\Delta}$-model, diminishing the benefits of self-ensembling. In contrast to using a fixed $\omega$, we adopt a decay function, where $\omega$ dynamically varies from 0.08 to 0.02 during training. This strategy achieves a good trade-off between the consistency and diversity of the perturbed models.
@@
To better understand the presented perturbation mechanism, in Fig. 11, we illustrate the percentage of perturbed Gaussians in the $\mathbf{\Delta}$-model. Note that the threshold $\tau$ in Eq. 15 is defined adaptively. Therefore, the percentage dynamically varies throughout training. This dynamic behavior balances the diversity and reliability of the perturbed models, thereby enhancing the self-ensembling process.
@@
We provide more visualizations including pseudo-view renderings of the $\mathbf{\Sigma}$-model and the $\mathbf{\Delta}$-model, uncertainty maps, and NVS results. For more details, please refer to the uploaded videos. nvs_res.mov: Shows the novel-view renderings of 3DGS, CoR-GS, and our SE-GS, where our method results in finer details and fewer artifacts. uncertainty_map_update.mov: Records the update of uncertainty maps during training. For each scene, we sample a pseudo view and store the corresponding rendering of the $\mathbf{\Delta}$-model in the image buffer. The uncertainty map is then derived from this buffer. The updates across different training iterations are shown in the video. sigma_model_vs_perturbed_model.mov: Displays the pseudo-view renderings of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model used in the regularization Eq. 13.